OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 @agent0smith

OpenStudy (agent0smith):

If the 14th term is -33, and the 15th term is 9, the common difference must be 33+9 so 42. use \[\Large a_n = a_1 + d(n-1)\]and a15=9 so: \[\Large 9 = a_1 + 42(15-1)\]solve for a1

OpenStudy (anonymous):

an = -579 + 42(n - 1) @agent0smith

OpenStudy (agent0smith):

Yep

OpenStudy (anonymous):

YAY

OpenStudy (agent0smith):

:D